# Topological Data Analysis for Neural Network Analysis: A Comprehensive Survey

### 10 Dec 2023 v1

@@ 2.3.1 Persistent homology
Manifolds also yield persistence diagrams computed from Morse functions. A Morse function is a smooth function $f\colon\mathbb{M}\to\mathbb{R}$ on a smooth manifold $M$ such that all critical points are non-degenerate, that is, the Hessian of $f$ at each critical point is non-singular. Homology is not only defined for simplicial complexes, but for arbitrary topological spaces. Similarly to the simplicial case, given $n\in\mathbb{N}$ and a field $\mathbb{F}$, one can build persistence modules from the homology of sublevel sets of a Morse function at different levels $t\in\mathbb{R}$. Thus, the persistence module $\mathbb{V}_{n}^{\mathbb{F}}(f)$ for a Morse function $f$ on a smooth manifold is defined as
$V_{n}^{\mathbb{F}}(f)_{t}=H_{n}(f^{-1}(-\infty,t])),$
where homology is meant with coefficients in $\mathbb{F}$. If $M$ is compact, then $\mathbb{V}_{n}^{\mathbb{F}}(f)$ is of finite type and hence a barcode and a persistence diagram can be extracted from it.
@@
$W_{\infty}\left(D_{1},D_{2}\right)=\inf_{\eta\colon D_{1}^{\Delta}\to D_{2}^{\Delta}}\;\sup_{x\in D_{1}^{\Delta}}\lVert x-\eta(x)\rVert_{\infty},$
where $D_{1}^{\Delta}$ and $D_{2}^{\Delta}$ denote the persistence diagrams $D_{1}$ and $D_{2}$ supplemented with their diagonals, that is, $D_{1}^{\Delta}=D_{1}\cup\Delta$ and $D_{2}^{\Delta}=D_{2}\cup\Delta$, where
$\Delta=\bigcup_{i=1}^{\infty}\left\{(x,x):x\in\bar{\mathbb{R}}\right\},$
and $\eta\colon D_{1}^{\Delta}\to D_{2}^{\Delta}$ are bijections of multisets, where points in $\Delta$ have countably infinite multiplicity. The Wasserstein distance $W_{q}$ is defined as
$W_{q}(D_{1},D_{2})=\inf_{\eta\colon D_{1}^{\Delta}\to D_{2}^{\Delta}}\Bigg{(}\,\sum_{x\in D_{1}^{\Delta}}\lVert x-\eta(x)\rVert_{\infty}^{q}\Bigg{)}^{1/q}$
@@
2.3.2 Mapper
@@
Mapper is an algorithm, introduced by Singh et al. (2007), that extracts visual descriptions of high-dimensional datasets in the form of simplicial complexes (usually graphs). One key concept behind Mapper is the nerve of a covering. Given a topological space $X$ and a covering $\mathcal{U}=\left\{\mathcal{U}_{i}\right\}_{i\in I}$ of $X$, the nerve of $\mathcal{U}$ is defined as the simplicial complex $N\left(\mathcal{U}\right)$ whose set of vertices is $I$ and where a family $\left\{i_{1},\ldots,i_{k}\right\}\subseteq I$ is a simplex of $N\left(\mathcal{U}\right)$ if and only if $\bigcap_{j=1}^{k}\mathcal{U}_{i_{j}}\neq\emptyset$. For well-behaved coverings, the geometric realization of the nerve $\mathcal{N}\left(\mathcal{U}\right)$ is homotopy equivalent to the space $X$. Therefore, $\mathcal{N}\left(\mathcal{U}\right)$ encodes relevant features of the topology of $X$ into a combinatorial object.
Mapper applies the nerve construction to suitably customized coverings of datasets. Simplicial complexes produced by Mapper are often able to reveal patterns of a latent topological space from which the dataset might have been sampled.
Mapper has the following ingredients as input: 1. A dataset $\mathcal{D}$, which we assume of finite cardinality $m$; 2. A filter function $f\colon\mathcal{D}\to\mathbb{R}^{d}$, where $d=1$ or $d=2$ in most use cases; 3. A finite cover $\mathcal{U}=\left\{\mathcal{U}_{i}\right\}_{i\in I}$ of $\text{Im}(f)\subset\mathbb{R}^{d}$ consisting of overlapping open sets; 4. A clustering algorithm, such as DBSCAN (Ester et al., 1996). The Mapper algorithm is described below in Algorithm 1.
Algorithm 1 Mapper algorithm
0:  Simplicial complex $S_{\mathcal{D}}$.
1:  $S_{\mathcal{D}}\leftarrow\emptyset$
3:  for all $i\in I$ do
6:  end for
8:     if $\bigcap_{j=1}^{t}C_{j}\neq\emptyset$ then
10:     end if
11:  end for
12:  return  $S_{\mathcal{D}}$
Mapper is most often applied using filter functions $f\colon\mathcal{D}\to\mathbb{R}$ and choosing covers $\mathcal{U}$ consisting of intervals with a fixed percentage of overlap $p$. With these choices and using clustering algorithms that generate disjoint clusters, the simplicial complexes produced by Mapper are graphs. Mapper graphs are closely related with Reeb graphs (Reeb, 1946). For an advanced analysis of the Mapper algorithm for filter functions $f\colon\mathcal{D}\to\mathbb{R}$ and its connection with Reeb graphs, we refer the reader to the work by Carrière and Oudot (2018).
@@ 2.4.1 Regularization
In machine learning, regularization techniques are algorithmic tweaks intended to reward models of lower complexity (Zhang et al., 2021). In the context of deep learning, regularization techniques aim to reduce model overfitting, that is, to avoid neural networks that excel in minimizing empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}$ but generalize poorly to a low real risk $\mathcal{R}$.
@@
Regularization terms are (almost everywhere) differentiable functions $\mathcal{T}\colon\Theta\to\mathbb{R}$ depending on the parameters of the neural network architecture being trained. They are minimized next to the empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$, i.e., making the training algorithm minimize the quantity
^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)+\alpha\mathcal{T}(\theta),$
@@ 2.4.6 Prediction of accuracy
Assuming that the values of $\mathcal{Y}$ are fully determined by a function $f\colon\mathcal{X}\to\mathcal{Y}$ in the data distribution, the real risk $\mathcal{R}(\mathcal{N})$ for this loss function is the expectation that the neural network $\mathcal{N}$ does not match the function $f$ for an input value $x$, that is, missclassifies $x$. If instead we consider the value $1-\mathcal{R}(\mathcal{N})$, this is the expectation that the neural network correctly classifies an input $x$, i.e., the probability of $\mathcal{N}$ matching $f$. We define this value by accuracy and denote it by $\text{Acc}(\mathcal{N})=1-\mathcal{R}(\mathcal{N})$. The higher the accuracy, the better the model according to the risk of $\mathcal{L}$. The problem with $\text{Acc}(\mathcal{N})$ is that we cannot compute it, so we approximate it by the empirical accuracy of $\mathcal{N}$ in a sample $\mathcal{D}$ different from the training dataset $\mathcal{D}_{\text{train}}$, usually called test dataset and denoted by $\mathcal{D}_{\text{test}}$. Therefore, this empirical accuracy in $\mathcal{D}$ can be computed as
@@ 2.4.7 Quality assessment of generative models
Neural network quality measurements in classification tasks are more straightforward to define than in generative tasks. In general, there are many interesting factors that can be evaluated to compare the quality of two generative neural networks, and thus deciding which is better is a hard challenge. For example, one of these factors could be the similarity between the supports of the original data distribution $\mathbb{P}_{X}$ and the distribution induced by the generation function $G$. Another interesting factor is the disentanglement capacity of the model $G$, which measures the ability of a model to separate (in a known way) the underlying factors of variation in the data it generates.
@@ 3 Analysis of neural networks using topological data analysis
3.1 Structure of the neural network
@@
Chowdhury et al. (2019) studied and compared two special homology groups designed for directed graphs in the context of FCFNNs. More specifically, the ranks of these two variants of homology groups are computed. These groups are the path homology (Grigor’yan et al., 2014) and the directed flag complex (DFC) homology (Reimann et al., 2017) groups, denoted by $\text{PathHom}_{k}(G(\mathcal{N}))$ and $\text{DFCHom}_{k}(G(\mathcal{N}))$, respectively, where $k$ denotes the dimension of the homology groups as in the simplicial case. In particular, Chowdhury et al. showed that, given a neural network $\mathcal{N}$ with architecture $a=(N,\varphi)$ and $N=(N_{0},\ldots,N_{L})$, and dimension $k\in\mathbb{N}$, the following holds:
@@
where $\text{rank}\left(H_{k}(G(\mathcal{N}))\right)$ is the rank of the $k$-dimensional simplicial homology group of $G(\mathcal{N})$ seen as an undirected graph. It can be further shown that these ranks are given by
@@ 3.2 Input and output spaces
In this section, we analyze how TDA has been used to analyze decision regions, boundaries, decompositions of the input space, and input and output spaces of generative models.
Decision regions
@@
A problem with the previous constructions is that they may require very large samples to robustly recover the topology of the decision boundaries. Obtaining such samples can sometimes be infeasible, especially when the cost of evaluating the value of $y\in\mathcal{Y}$ by the unknown function $f$ given $x\in\mathcal{X}$ is high. Li et al. (2020) propose a method based on active learning (Settles, 2009) to sample a small but meaningful set to recover the homology of the decision boundaries using the labeled Čech and Vietoris–Rips complexes introduced in the previous article. Information extracted from persistence diagrams is used in both articles to perform model selection for basic datasets including Banknote (Lohweg, 2013), CIFAR-10 (Krizhevsky, 2009), MNIST (Deng, 2012), and Fashion-MNIST (Xiao et al., 2017), and basic machine and deep learning models.
@@
Neural networks with a ReLU activation function $\varphi(x)=\max(0,x)$ determine a decomposition of the input space into convex polyhedra that assigns to each polyhedron in the decomposition a unique binary vector in such a way that two polyhedra share a facet if and only if their associated binary vectors differ exactly in one component. Liu et al. (2023) use this polyhedral decomposition of the input space to infer the topology of manifolds embedded in the domain of a given ReLU neural network from a sample of their points using persistence diagrams. To compute persistence diagrams given the sample $S=\{x_{i}\}_{i=1}^{m}$ of the manifold $\mathbb{M}$, the proposed procedure computes the set of unique binary vectors $\mathfrak{B}(S)=\left\{b(x_{i})\right\}_{i=1}^{m}$ associated with the points of the sample, where $b(x_{i})$ is the binary vector associated with the convex polyhedra in which $x_{i}$ lies, and then computes the Vietoris–Rips persistence diagrams $D(\mathbb{V}_{k}(\text{VR}(\mathfrak{B}(S),h)))$ of $\mathfrak{B}(S)$ using the Hamming distance $h$ which counts the number of different bits in the two binary strings to compare.
Although this is not directly related to neural network analysis, we firmly believe that this opens the way to study the structure and topology of the unknown data distribution for a specific learning problem by applying this method to large enough data samples.
@@
In a very similar fashion, but with a simpler approach, Charlier et al. (2019) propose to compare generative models $\mathcal{G}$, this time not necessarily restricted to GAN models, by directly comparing the zero and one-dimensional persistence diagrams induced by Vietoris–Rips filtrations of samples from $\mathcal{X}$ and $\text{Im}\left(\phi_{\mathcal{G}}\right)$ distributed by $\mathbb{P}_{X}$ and $\phi_{\mathcal{G}}\left(\mathbb{P}_{Z}\right)$ respectively, using the Bottleneck distance. In this case, the experiments compared four different generative models in the credit card fraud detection dataset (cre, ). These models were WGAN and WGAN-GP together with WAE (Tolstikhin et al., 2018) and VAE (Kingma and Welling, 2014). In this case too, WGAN-GP was shown to produce better results than the other models.
@@
Generative models aim to control the attributes of the generated samples. In particular, a key goal is to design generative models $\mathcal{G}$ that map from a latent space $\mathcal{Z}$ to a data space $\mathcal{Y}$ via $\phi_{\mathcal{G}}$ in such a way that the latent space $\mathcal{Z}$ can be factored into subspaces corresponding to factors of variation of the data generated in such a way that, if we change the value of a subspace associated with a factor of variation in the latent space, the generated data change only in this source of variation. A model that satisfies this property is said to be disentangled. For a formal discussion of disentanglement, we refer the reader to Higgins et al. (2018). Measuring a model’s entanglement is challenging, and there is no canonical way to do it. Zhou et al. (2021) propose two measures of disentanglement, one unsupervised and one supervised, using topological data analysis. Although they perform similarly to other disentanglement metrics, these measures do not require some strong assumptions about the dataset or the model, except for the (weak) hypothesis that $\text{Im}(\phi_{\mathcal{G}})$ is a manifold for the unsupervised case and the more unrealistic assumption that the data space is a manifold for the supervised case. Roughly, these measures are based on two fundamental ideas for disentangled models $\mathcal{G}$: 1. The family of submanifolds of $\text{Im}(\phi_{\mathcal{G}})$ yielded by the images of $\phi_{\mathcal{G}}$ restricted to the different values of the same subspace corresponding to a factor of variation are pairwise homeomorphic and thus have the same topology; 2. Submanifolds corresponding to different factors of variation are usually not homeomorphic.

### 03 Jan 2024 v2

@@ 2.3.1 Persistent homology
Manifolds also yield persistence diagrams computed from Morse functions. A Morse function is a smooth function $f\colon\mathbb{M}\to\mathbb{R}$ on a smooth manifold $M$ such that all critical points are non-degenerate, that is, the Hessian of $f$ at each critical point is non-singular. Similarly to the simplicial case, given $n\in\mathbb{N}$ and a field $\mathbb{F}$, one can build persistence modules from the homology of sublevel sets of a Morse function at different levels $t\in\mathbb{R}$. Thus, the persistence module $\mathbb{V}_{n}^{\mathbb{F}}(f)$ for a Morse function $f$ on a smooth manifold is defined as
$V_{n}^{\mathbb{F}}(f)_{t}=H_{n}(f^{-1}(-\infty,t]),$
where singular homology is meant with coefficients in $\mathbb{F}$. If $M$ is compact, then $\mathbb{V}_{n}^{\mathbb{F}}(f)$ is of finite type and hence a barcode and a persistence diagram can be extracted from it.
@@
$W_{\infty}\left(D_{1},D_{2}\right)=\inf_{\eta:D_{1}^{\Delta}\to D_{2}^{\Delta}}\;\sup_{x\in D_{1}^{\Delta}}\lVert x-\eta(x)\rVert_{\infty},$
where $D_{1}^{\Delta}$ and $D_{2}^{\Delta}$ denote the persistence diagrams $D_{1}$ and $D_{2}$ supplemented with their diagonals, that is, $D_{1}^{\Delta}=D_{1}\cup\Delta$ and $D_{2}^{\Delta}=D_{2}\cup\Delta$, where $\Delta=\left\{(x,x):x\in\bar{\mathbb{R}}\right\}$, and $\eta\colon D_{1}^{\Delta}\to D_{2}^{\Delta}$ are bijections of multisets, where points in $\Delta$ have countably infinite multiplicity. The Wasserstein distance $W_{q}$ is defined as
$W_{q}(D_{1},D_{2})=\inf_{\eta:D_{1}^{\Delta}\to D_{2}^{\Delta}}\Bigg{(}\,\sum_{x\in D_{1}^{\Delta}}\lVert x-\eta(x)\rVert_{\infty}^{q}\Bigg{)}^{1/q}$
@@
2.3.2 Mapper and GTDA
Mapper is an algorithm, introduced by Singh et al. (2007), that extracts visual descriptions of high-dimensional datasets in the form of simplicial complexes (usually graphs). One key concept behind Mapper is the nerve of a covering. Given a topological space $X$ and a covering $\mathcal{U}=\left\{\mathcal{U}_{i}\right\}_{i\in I}$ of $X$, the nerve of $\mathcal{U}$ is defined as the simplicial complex $N\left(\mathcal{U}\right)$ whose set of vertices is $I$ and where a family $\left\{i_{1},\ldots,i_{k}\right\}\subseteq I$ is a simplex of $N\left(\mathcal{U}\right)$ if and only if $\bigcap_{j=1}^{k}\mathcal{U}_{i_{j}}\neq\emptyset$. For good coverings (open coverings whose sets and their non-empty finite intersections are contractible), the geometric realization of $\mathcal{N}\left(\mathcal{U}\right)$ is homotopy equivalent to $X$. Hence, $\mathcal{N}\left(\mathcal{U}\right)$ encodes relevant features of the shape of $X$ into a combinatorial object.
@@ 2.3.2 Mapper and GTDA
Mapper applies the nerve construction to suitably customized coverings of datasets. Simplicial complexes produced by Mapper are often able to reveal patterns of a latent topological space from which the dataset might have been sampled. The Mapper algorithm has the following ingredients as input: 1. A dataset $\mathcal{D}$, which we assume of finite cardinality; 2. A filter function $f\colon\mathcal{D}\to\mathbb{R}^{d}$, where $d=1$ or $d=2$ in most use cases; 3. A finite cover $\mathcal{U}=\left\{\mathcal{U}_{i}\right\}_{i\in I}$ of $\text{Im}(f)\subset\mathbb{R}^{d}$ consisting of overlapping open sets; 4. A clustering algorithm, such as DBSCAN (Ester et al., 1996). The method is described in Algorithm 1 in the Appendix.
Mapper is most often applied using filter functions $f\colon\mathcal{D}\to\mathbb{R}$ and choosing covers $\mathcal{U}$ consisting of intervals with a fixed percentage of overlap $p$. With these choices and using clustering algorithms that generate disjoint clusters, the simplicial complexes produced by Mapper are graphs. Mapper graphs are closely related with Reeb graphs (Reeb, 1946; Edelsbrunner and Harer, 2022, Chapter VI). For an advanced analysis of the Mapper algorithm for filter functions $f\colon\mathcal{D}\to\mathbb{R}$ and its connection with Reeb graphs, we refer the reader to the work by Carrière and Oudot (2018).
It was observed by Liu et al. (2023a) that, in some experiments studying the output space of neural networks, Mapper produced too many tiny components that did not carry useful information for the experiments. For this reason, they introduced a new algorithm called graph-based topological data analysis (GTDA). The GTDA algorithm builds on the Mapper algorithm to construct Reeb networks from graph inputs instead of point clouds. Reeb networks generalize Reeb graphs. Each vertex in a Reeb network built using GTDA is associated with a subgraph of the original graph. Given an input graph $G$, GTDA uses filter functions $f\colon V(G)\to\mathbb{R}^{d}$ to build Reeb network vertices using a recursive splitting strategy based on filter function values. After initial vertices are generated, the smallest ones are merged into other vertices and edges are added. Vertices within connected components with non-empty intersection are connected, and other edges are added to promote connectivity of the output. The GTDA algorithm is described in Liu et al. (2023a, Algorithm 1).
@@ 2.4.1 Regularization
In machine learning, regularization techniques are algorithmic tweaks intended to reward models of lower complexity (Zhang et al., 2021). In the context of deep learning, regularization techniques aim to reduce model overfitting, that is, to avoid neural networks that excel in minimizing empirical risk ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}$ but generalize poorly to a low real risk $\mathcal{R}$.
@@
Regularization terms are (almost everywhere) differentiable functions $\mathcal{T}\colon\Theta\to\mathbb{R}$ depending on the parameters of the neural network architecture being trained. They are minimized next to the empirical risk ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$, i.e., making the training algorithm minimize the quantity
^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)+\alpha\mathcal{T}(\theta),$
@@ 2.4.6 Prediction of accuracy
Assuming that the values of $\mathcal{Y}$ are fully determined by a function $f\colon\mathcal{X}\to\mathcal{Y}$ in the data distribution, the real risk $\mathcal{R}(\mathcal{N})$ for this loss function is the expectation that the neural network $\mathcal{N}$ does not match the function $f$ for an input value $x$, that is, misclassifies $x$. If instead we consider the value $1-\mathcal{R}(\mathcal{N})$, this is the expectation that the neural network correctly classifies an input $x$, i.e., the probability of $\mathcal{N}$ matching $f$. We define this value by accuracy and denote it by $\text{Acc}(\mathcal{N})=1-\mathcal{R}(\mathcal{N})$. The higher the accuracy, the better the model according to the risk of $\mathcal{L}$. The problem with $\text{Acc}(\mathcal{N})$ is that we cannot compute it, so we approximate it by the empirical accuracy of $\mathcal{N}$ in a sample $\mathcal{D}$ different from the training dataset $\mathcal{D}_{\text{train}}$, usually called test dataset and denoted by $\mathcal{D}_{\text{test}}$. Therefore, this empirical accuracy in $\mathcal{D}$ can be computed as
@@ 2.4.7 Quality assessment of generative models
Neural network quality measurements in classification tasks are more straightforward to define than in generative tasks. In general, there are many interesting factors that can be evaluated to compare the quality of two generative neural networks, and thus deciding which is better is a hard challenge. For example, one of these factors could be the similarity between the supports of the original data distribution $\mathbb{P}_{X}$ and the distribution induced by the generation function.
@@ 3 Analysis of neural networks using topological data analysis
3.1 Structure of a neural network
@@
Chowdhury et al. (2019) studied and compared two special homology groups designed for directed graphs in the context of FCFNNs. More specifically, the ranks of these two variants of homology groups are computed. These groups are the path homology (Grigor’yan et al., 2014) and the directed flag complex (DFC) homology (Reimann et al., 2017) groups, denoted by $\text{PathHom}_{k}(G(\mathcal{N}))$ and $\text{DFCHom}_{k}(G(\mathcal{N}))$, respectively, where $k$ denotes the dimension of the homology groups as in the simplicial case. In particular, Chowdhury et al. showed that, given a neural network $\mathcal{N}$ with architecture $a=(N,\varphi)$, where $N=(N_{0},\ldots,N_{L})$, and given a dimension $k\in\mathbb{N}$, the following equalities hold:
@@
where $\text{rank}\left(H_{k}(G(\mathcal{N}))\right)$ is the rank of the $k$-dimensional simplicial homology group of $G(\mathcal{N})$ seen as an undirected graph. It can be further shown that
@@ 3.2 Input and output spaces
In this section, we analyze how TDA has been used to analyze decision regions, boundaries, decompositions of the input space, and input and output spaces of generative and non-generative models.
General output space and decision regions
@@
On a more general setting, Liu et al. (2023a) proposed the GTDA algorithm of Section 2.3.2 as an improved version of Mapper to study the output space of general neural network functions. They studied Reeb networks built using GTDA with input graphs induced from the output vectors $(\phi_{\mathcal{N}}^{(L)}(x)_{1},\ldots,\phi_{\mathcal{N}}^{(L)}(x)_{N_{L}})$ of neural network functions given a dataset $\mathcal{D}$ and filter functions given by the same output values plus, possibly, some extra values depending on the dataset or the outputs. To give output values a graph structure, they either used previously-known binary relationships between the data or binary relationships given by nearest neighbors algorithms as edges.
Liu et al. were able to gain useful information from GTDA graphs that was not found using Mapper or other alternatives about output values of the Enformer model (Avsec et al., 2021) used as inputs to predict harmful mutations of a human gene. Also, GTDA graphs were used to perform automatic error estimation for machine learning prediction tasks. A numerical error estimation score for a sample $x$ was calculated by comparing the predicted label of $x$ with ground truth labels of training and validation samples whose prediction vectors were surrounding the prediction vector of $x$ in the union of subgraphs of the Reeb network vertices plus the extra edges added in the GTDA algorithm. The error estimation score was shown to successfully correct erroneous predictions in binary classification tasks. Also, it detected regions of the output space with many misclassifications, that allowed them to determine the source of error in those regions for the previous classification problem dealing with the Enformer architecture. Similar experiments were performed on other network architectures and datasets such as ResNet-50 and Imagenette (Howard, 2021), showing that GTDA offers unique insights about the output space that cannot be recovered with other methods such as Mapper.
@@
A problem with the previous constructions is that they may require very large samples to robustly recover the topology of the decision boundaries. Obtaining such samples can sometimes be infeasible, especially when the cost of evaluating the value of $y\in\mathcal{Y}$ by the unknown function $f$ given $x\in\mathcal{X}$ is high. Li et al. (2020) propose a method based on active learning (Settles, 2009) to sample a small but meaningful set to recover the homology of the decision boundaries using the labeled Čech and Vietoris–Rips complexes introduced in the article. Information extracted from persistence diagrams is used in both articles to perform model selection for basic datasets including Banknote (Lohweg, 2013), CIFAR-10 (Krizhevsky, 2009), MNIST (Deng, 2012), and Fashion-MNIST (Xiao et al., 2017), and basic machine and deep learning models.
@@
Neural networks with a ReLU activation function $\varphi(x)=\max(0,x)$ determine a decomposition of the input space into convex polyhedra that assigns to each polyhedron in the decomposition a unique binary vector in such a way that two polyhedra share a facet if and only if their associated binary vectors differ exactly in one component. Liu et al. (2023b) use this polyhedral decomposition of the input space to infer the topology of manifolds embedded in the domain of a given ReLU neural network from a sample of their points using persistence diagrams. To compute persistence diagrams given the sample $S=\{x_{i}\}_{i=1}^{m}$ of the manifold $\mathbb{M}$, the proposed procedure computes the set of unique binary vectors $\mathfrak{B}(S)=\left\{b(x_{i})\right\}_{i=1}^{m}$ associated with the points of the sample, where $b(x_{i})$ is the binary vector associated with the convex polyhedra in which $x_{i}$ lies, and then computes the Vietoris–Rips persistence diagrams $D(\mathbb{V}_{k}(\text{VR}(\mathfrak{B}(S),h)))$ of $\mathfrak{B}(S)$ using the Hamming distance $h$, which counts the number of different bits in the two binary strings to compare. Although this is not directly related to neural network analysis, we firmly believe that this opens a way to study the structure and topology of the unknown data distribution for a specific learning problem by applying this method to large enough data samples.
Masden (2022) used polyhedral decompositions to study the topology of decision boundaries of small, non-trained, ReLU neural networks by attaching a polyhedral complex structure to the decomposition. Polyhedral complexes are similar to simplicial complexes, in that they are sets of polyhedra such that every face of a polyhedron of the complex is also in the complex, and the intersection of any two polyhedra is either empty or a face of both. Masden found that polyhedral complexes derived from neural networks under some technical assumptions possess a dual structure known as signed sequence cubical complexes, and provided a polynomial time algorithm with respect to the number of non-input neurons to compute both complexes, yet exponential with respect to the input dimension. The decision boundary of the neural network corresponds to specific polyhedra, which in turn have dual representations in the signed sequence cubical complexes, from which homology groups for the decision boundary can be recovered under certain assumptions, such as the need of performing a one-point compactification of the decision boundary.
Differences were found in the structure of decision boundaries in terms of their neural network layer structures. In networks with a single hidden layer, the topology of decision boundaries exhibited remarkable consistency across various parameter ranges, including modifications to input data dimension and the number of hidden neurons. In contrast, networks with two hidden layers displayed more variability in the topological distribution of their decision boundaries.
@@
In a very similar fashion, but with a simpler approach, Charlier et al. (2019) propose to compare generative models, this time not necessarily restricted to GAN models, by directly comparing the zero and one-dimensional persistence diagrams induced by Vietoris–Rips filtrations of samples taken from the real data distribution and from the generative model using the bottleneck distance. In this case, the experiments compared four different generative models in the credit card fraud detection dataset (cre, ). These models were WGAN and WGAN-GP together with WAE (Tolstikhin et al., 2018) and VAE (Kingma and Welling, 2014). In this case too, WGAN-GP was shown to produce better results than the other models.
The two aforementioned approaches evaluating the quality of generative models were based on comparing the topology of real and synthetic data by means of persistent homology. However, they are not the usual choices in the generative deep learning community, where scores such as the Fréchet inception distance (Heusel et al., 2017) or a numerical approximation of precision and recall (Sajjadi et al., 2018) are currently preferred to evaluate generative models. Kim et al. (2023) propose an alternative way to approximate precision and recall scores using an approximation of the support of the data distributions based on preimages of kernel density estimators (KDE) that preserve the topology of the support of a smoothed version of the real and synthetic data distributions. Their approximation is based on results of Fasy et al. (2014, Method IV), where it is argued that persistence diagrams from superlevel set persistence modules may carry topological information from the support of the data distribution. Specifically, the supports of the real and synthetic data are approximated by the superlevel sets $(\widehat{p}_{h_{r}})^{-1}[c_{r},\infty)$ and $(\widehat{p}_{h_{s}})^{-1}[c_{s},\infty)$, respectively, where $\widehat{p}_{h_{r}}$ and $\widehat{p}_{h_{s}}$ are KDEs for real and synthetic datasets $\mathcal{D}_{r}$ and $\mathcal{D}_{s}$ given by
$\widehat{p}_{h_{\bullet}}(x)=\frac{1}{\left|\mathcal{D_{\bullet}}\right|}\sum_{p\in\mathcal{D}_{\bullet}}\frac{1}{h^{d}}\,K\left(\frac{x-p}{h}\right),$
where $h>0$ and $K$ are the bandwidth and kernel of the KDEs, selected beforehand, and $c_{r}$ and $c_{s}$ are confidence bands for a given significance value $\alpha$, obtained using bootstrap over the datasets $\mathcal{D}_{r}$ and $\mathcal{D}_{s}$ with the condition that
$\liminf_{\left|\mathcal{D}_{\bullet}\right|\to\infty}\;\mathbb{P}\left(\lVert\widehat{p}_{h_{\bullet}}-p_{h_{\bullet}}\rVert_{\infty}<c_{\bullet}\right)\geq 1-\alpha,$
where $p_{h_{r}}$ and $p_{h_{s}}$ are smoothed versions of the real and synthetic data distributions, respectively (Fasy et al., 2014, Equation (26)). By the stability theorems (Cohen-Steiner et al., 2007), the confidence bands $c_{r}$ and $c_{s}$, also bound the distances between the persistence diagrams coming from the functions $\widehat{p}_{h_{\bullet}}$ and $p_{h_{\bullet}}$ with probability $1-\alpha$, allowing one to study which points in the persistence diagrams generated from the data samples $\mathcal{D}_{\bullet}$ by $\widehat{p}_{h_{\bullet}}$ can be considered noise with respect to the persistence diagrams generated by $p_{h_{\bullet}}$; see Fasy et al. (2014, Section 4). The superlevel sets $(\widehat{p}_{h_{\bullet}})^{-1}[c_{\bullet},\infty)$ are regions which intend to induce persistence diagrams without noisy points according to the previous confidence bands, thus recovering supports with similar topology to the topology of the support of the smoothed data distributions measured by the distances between their respective persistence diagrams.
Under some technical assumptions, the precision and recall scores approximated using the superlevel sets proposed by Kim et al. become close to the real precision and recall from the distribution as more examples are added to the datasets $\mathcal{D}_{r}$ and $\mathcal{D}_{s}$. The approximation is done with robustness, meaning that it holds even with data possibly corrupted by noise. The suitability of the new metrics was tested on several synthetic and real scenarios, where the metric accurately describes the differences between real and synthetic data distributions in several scenarios where other metrics struggle to evaluate differences. Also, the F1-score computed from the precision and recall approximations ranked different generative models such as StyleGAN2 (Karras et al., 2020), ReACGAN (Kang et al., 2021), among others, as the FID metric, the primary metric to score generative models in the moment of the publication of this survey, making this metric consistent with the state-of-the-art knowledge about the quality of generative models.
A class of generative models create synthetic data transforming values from a latent space $\mathcal{Z}$ to the real data space $\mathcal{X}$ via a function $G\colon\mathcal{Z}\to\mathcal{X}$. For such generative models, a desirable property is to be able to control the attributes of the generated samples. One approach to controlling attributes involves designing generative models that map from the latent space to the data space in such a way that the latent space $\mathcal{Z}$ can be factored into subspaces corresponding to factors of variation of the data generated, meaning that if one changes the value of a subspace associated with a factor of variation in the latent space, then the generated data change only in this source of variation. A model that satisfies this property is said to be disentangled. For a formal discussion of disentanglement, we refer the reader to Higgins et al. (2018). Preliminary work on disentangling generative models using topological data analysis to build regularization terms can be found in Balabin et al. (2023). In this work, regularization terms are based on the differences between generated data from two samples from the latent space using the representation topology divergence, a method to compare point clouds topologically which is reviewed in Section 3.3.2.
@@
Measuring a model’s entanglement is challenging, and there is no canonical way to do it. Zhou et al. (2021) propose two measures of disentanglement, one unsupervised and one supervised, using topological data analysis. Although they perform similarly to other disentanglement metrics, these measures do not require some strong assumptions about the dataset or the model, except for the (weak) hypothesis that $\text{Im}(G)$ is a manifold for the unsupervised case and the more unrealistic assumption that the data space is a manifold for the supervised case. Roughly, these measures are based on two fundamental ideas for disentangled models: 1. The family of submanifolds of $\text{Im}(G)$ yielded by the images of $G$ restricted to the different values of the same subspace corresponding to a factor of variation are pairwise homeomorphic and thus have the same topology; 2. Submanifolds corresponding to different factors of variation are usually not homeomorphic.